a carrom board (4X4 ft sqr.) has the queen at the centre. the queen , hit by the stricker moves to the front edge, rebounds and goes in the hole behind the stricking line . find the magnitude of displacement of the queen (a) from the cenre to front edge (b) from the front edge to the hole (c) from the centre to the holom the front edge to the hole (c) from the centre to the hole

Grade:11

1 Answers

Vikas TU
3 years ago
In triangle ABC,
tan Ф = x/2 ................. equation 1
In triangle DCE,
tan Ф = (2-x)/4 ................... equation 2
Equating both equation 1 and 2
x/2 = (2-x)/4
2x = 2-x
3x = 2
x = 2/3 ft.
(a) In triangle ABC (by Pythagoras theorem)
      AC = √(AB² + BC²) = (2√10)/3 ft.
(b) In triangle CDE ,
     DE = 1 - (2/3) = 4/3 ft.
     CD = 4ft.
So, by Pythagoras Theorem 
     CE = √(CD² + DE²) = (4√10)/3 ft.
(c) In triangle AGE , (by Pythagoras Theorem)
      AE = √(AG² + GE²) = 2√2 ft.